You have 35 years left until retirement and want to retire with $1.3 million. Your salary is paid annually, and you will receive $63,000 at the end of the current year. Your salary will increase at 2 percent per year, and you can earn a return of 8 percent on the money you invest. If you save a constant percentage of your salary, what percentage of your salary must you save each year?
Answer #1

Amount required at the time of retirement = $1.3 million (this is future value i.e. value at the end of 35th year)

Let the % saved of salary be x% each year. Thus amount saved at the end of current year = 63,000*x/100. Future value of this amount = 63,000*x/100*1.08^(35-1)

Salary in year 2 = 63,000*1.02 = 64,260. Future value of this amount = 64,260*x/100*1.08^(35-2)

Thus 63,000*x/100*1.08^(35-1)+63,000*1.02^1*x/100*1.08^(35-2).......+63,000*1.02^34*x/100*1.08^(35-35) = $1.3 million

Solving the above equation we get x = 9.6836% (or 9.68% after rounding off to 2 decimal place)

Year Salary % saved Savings Future value
1 63,000.00 9.68 6,100.68 83,519.16
2 64,260.00 6,222.70 78,879.21
3 65,545.20 6,347.15 74,497.03
4 66,856.10 6,474.09 70,358.30
5 68,193.23 6,603.57 66,449.51
6 69,557.09 6,735.65 62,757.87
7 70,948.23 6,870.36 59,271.32
8 72,367.20 7,007.77 55,978.47
9 73,814.54 7,147.92 52,868.55
10 75,290.83 7,290.88 49,931.41
11 76,796.65 7,436.70 47,157.45
12 78,332.58 7,585.43 44,537.59
13 79,899.23 7,737.14 42,063.28
14 81,497.22 7,891.88 39,726.43
15 83,127.16 8,049.72 37,519.40
16 84,789.71 8,210.72 35,434.99
17 86,485.50 8,374.93 33,466.38
18 88,215.21 8,542.43 31,607.14
19 89,979.51 8,713.28 29,851.19
20 91,779.10 8,887.54 28,192.79
21 93,614.69 9,065.29 26,626.52
22 95,486.98 9,246.60 25,147.27
23 97,396.72 9,431.53 23,750.20
24 99,344.65 9,620.16 22,430.74
25 101,331.55 9,812.56 21,184.59
26 103,358.18 10,008.82 20,007.67
27 105,425.34 10,208.99 18,896.13
28 107,533.85 10,413.17 17,846.35
29 109,684.52 10,621.44 16,854.88
30 111,878.22 10,833.86 15,918.50
31 114,115.78 11,050.54 15,034.14
32 116,398.10 11,271.55 14,198.91
33 118,726.06 11,496.98 13,410.08
34 121,100.58 11,726.92 12,665.08
35 123,522.59 11,961.46 11,961.46
Total 1,300,000.00